McDowell’s hamburger restaurant chain had revenues of $5,844,800 last year (year 1) and $7,504,500 this year.

1. What is the year on year sales growth? 28.4%
2. $445,593 of this year’s revenue came from the opening of a new location. What is your same stores sales growth? 20.77%
3. Why is it important to calculate same stores sales growth? Why not just use year on year instead?
4. Assume the company’s growth is the same for each year. What is the company’s projected sales in year 4?
5. Your year 6 actual sales volume is $10,545,700. What is your CAGR over the 6 years?

need help with the last three.
Answer #1

1. Sales growth= Sales of this year-Sales of last year/Sales of last year= 7504500-5844800/5844800= 28.4%

  • 2. Revenue from opening new location= 445593
  • Revenue in current year on same stores= 7504500-445593= 7058907
  • Growth rate= 7058907-5844800/5844800= 20.77%

3. It is important to calculate same years sales growth as it serves an important metrics for various decisions in the organisation. Even manager's bonus is linked to the sales growth. Sales growth rate gives an evidence if the sales goals are met which is not possible when year on year sales are analysed

4. Assuming company's growth is same for ech year i.e. 28.4%. Sales in Year 1= 5844800, sales in year 4= 5844800(1.284)^3= $12372707

5. CAGR= (Sales in year6/Sales in year 1)^1/t-1=(10545700/5844800)^1/6-1= 10.35% approx